A 60 g sample of iron reacts with 26 g of oxygen to form how many grams of iron oxide? Express your answer using two significant figures.
1
Expert's answer
2019-01-14T07:35:45-0500

4Fe + 3O2 = 2Fe2O3

n=m/M

n(Fe)= 60g/55.85g/mol = 1.074 mol

n(O2)= 26g / 32.00g/mol = 0.8125 mol

Find limiting reactant. Compare:

n(Fe)/4 and n(O2)/3

1.074/4 and 0.8125/3

0.2685<0.2708

So, Fe is a limiting reactant. Use amount of substance of Fe to find amount of substance of Fe2O3.

1.074/4 = n(Fe2O3)/2

n(Fe2O3)= 0.537 mol

m=M*n

M(Fe2O3) = 55.85*2+16.00*3= 159.7 g/mol

m(Fe2O3) = 0.537 mol *159.7g/mol = 85.76 g = 86g.

Answer: 86 g